OpenStudy (anonymous):

Solve this problem in your notebook using all four steps. Harvey is 3 times as old as Jane. The sum of their ages is 48 years. Find the age of each. Jane is______years old. Harvey is______years old.

OpenStudy (texaschic101):

write it like you read it... Harvey is 3 times as old as Jane... H = 3J The sum of their ages is 48 H + J = 48 so now we are just gonna sub in 3J for H back into the 2nd equation...solving by substitution. H + J = 48 3J + J = 48 4J = 48 -- divide both sides by 4 J = 12 H = 3J H = 3(12) H = 36 so...Harvey (H) is 36 and Jane (J) is 12